Show that if A, B, C, and D are sets with |A|=|B| and|C|=|D|, then |AxC|=|BxD|.
1
Expert's answer
2020-10-19T16:49:35-0400

The Cartesian product of two sets "X" and "Y", denoted "X\times Y", is the set of all ordered pairs "(x,y)" where "x" is in "X"and "y" is in "Y", that is "{\displaystyle X\times Y=\{\,(x,y)\mid x\in X\ {\text{ and }}\ y\in Y\,\}.}"

Since "|A|=|B|", there exists a bijection "f: A\to B". By analogy, the equality "|C|=|D|" implies that there exists a bijection "g: C\to D". Let us prove that the map "h: A\times C\to B\times D", "h(a,c)=(f(a), g(c))", is a bijection as well.

First prove that "h" is an injection. Let "h(a_1, c_1)=h(a_2, c_2)". So "(f(a_1), g(c_1))=(f(a_2), g(c_2))". It follows that "f(a_1)=f(a_2)" and "g(c_1)=g(c_2)". Since "f" and "g" are injections, "a_1=a_2" and "c_1=c_2". Therefore, "(a_1,c_1)=(a_2,c_2)", and "h" is an injection.

Finally, prove that "h" is a surjection. Let "(b,d)\in B\times D". Since "f" and "g" are surjections, there exist "a\in A" and "c\in C" such that "f(a)=b" and "g(c)=d". Then "h(a,c)=(f(a), g(c))=(b, d)", and consequantly, "h" is a surjection.

It follows that "h: A\times C\to B\times D" is bijection, and therefore, "|A\times C|=|B\times D|".